OpenStudy (anonymous):

find the average rate of change of f(x)=x^2+4 a) from 1 to 3 b) from 0 to 2 c) from -2 to 1

OpenStudy (anonymous):

sure branson

OpenStudy (anonymous):

dy/dx=2x

Parth (parthkohli):

Is it (x^2) + 4 or x^(2+4)

OpenStudy (anonymous):

2(3)-2(1)=4

OpenStudy (anonymous):

Two ways of doing it, plug in the limits of the interval to get y values, subtract them and divide by x (one end minus another). For instance: f(1) = 1^2 + 4 = 5 f(3) = 3^2 + 4 = 13 f(3) - f(1) = 8 3-1 = 2 Avg rate of change is 8/2 = 4.

OpenStudy (anonymous):

The other way is to take the derivative and take the average distance (x=2 for the first interval), as Tai is doing :-)

OpenStudy (anonymous):

2(2)-2(0)=4

OpenStudy (anonymous):

cheers bmp

OpenStudy (anonymous):

Ok.. I thought that when you do the rate of change it would look something like this.. f(x)=x^2 + 4 when imputing a) (from 1 to 3) the Awnser would be something like ... the average rate of change is from 5 - 13. I got that by... f(x)=(1)^2 + 4 = 5 f(x)=(3)^2 + 4 =13 so wouldnt the awnser to a be 5 to 13? Or am I doing it wrong?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok??? What do you mean? =) sorry Im just confused..lol

OpenStudy (anonymous):

No, you have to divide the x interval. You took sample points from 1 to 3, to it's 3-1. And similarly for the y-interval, you will get 13-5.

OpenStudy (anonymous):

Because you want the average, you are looking for a number, it's the average on a given interval. Remember:\[Slope = \frac{\Delta y}{\Delta x}\]This is the average rate of change. In general, you would take the derivative, but they are the same for the average case.

OpenStudy (anonymous):

Another way to think about it is to say "how fast is the function growing at a given point?" Your question is like that but with the change that you are asking is how fast the function is growing on average on a given interval. So, you want an average of the interval (halfway), that's why you have to subtract (or pick the midpoint for the derivative).

OpenStudy (anonymous):

So what is the actual awnser? It is giving me a place for 3 different awnsers...a,b,&c...hang on one second..I am trying tyo work it out

OpenStudy (anonymous):

For the first, it's 4. For the second, it's 2. For the last, it's 1.

OpenStudy (anonymous):

Typo, it's -1.

OpenStudy (anonymous):

yay...i checked my awnsers to your when I finished and I did it. Thank you so much! I wish you were my math teacher! =)

OpenStudy (anonymous):

You are welcome, but I am not good at teaching, haha.

OpenStudy (anonymous):

I can't tell hun!